### v1

@@ 3 Experimental Results
BEIR Zero-Shot Evaluations The BEIR (Benchmarking IR) benchmark is a comprehensive evaluation suite that encompasses a wide variety of Information Retrieval (IR) datasets designed to rigorously test the zero-shot performance of retrieval models. To comprehensively evaluate the performance of DeeperImpact  we benchmarked it against SPLADEv3-doc, SPLADEv3, and the large efficient SPLADE variant BT-SPLADE-L [40] on a subset of BEIR datasets using NDCG@10 as the metric. The zero-shot inferences were performed with DeeperImpact model trained on MSMARCO passage dataset. We also used the Llama 2 model—fine-tuned for query prediction on MSMARCO passage dataset—for expanding each of these datasets with relevant queries. The results are displayed in Table 5.
Among the SPLADE variants, SPLADE-v3 consistently achieves the highest NDCG@10 scores across the datasets, demonstrating its state-of-the-art performance. However, it is worth noting that SPLADEv3 incorporates query processing, making direct comparisons less straightforward.
SPLADE-v3-doc, which processes queries as a bag of words without additional query processing, is a more appropriate baseline for comparison with DeeperImpact. In this context, DeeperImpact exhibits competitive performance, often closely aligning with SPLADEv3-doc. Its performance is also fairly close with the large efficient SPLADE variant.
@@
Table 5: NDCG@10 Results on a Subset of BEIR Datasets
@@ 4 Conclusions and Future Work
In this paper, we presented DeeperImpact  an optimized version of the DeepImpact sparse retrieval model, aimed at improving retrieval effectiveness while maintaining computational efficiency. By substituting DocT5Query with a fine-tuned Llama 2 model for document expansion, leveraging advanced training strategies such as initializing with CoCondenser, employing hard negatives, and integrating distillation techniques, DeeperImpact significantly narrows the effectiveness gap with state-of-the-art but more expensive sparse retrieval models such as SPLADE. Future work could focus on further enhancing the training of DeeperImpact using a combination of training loss objectives and searching for optimal hyperparameters similar to what has been done with recent versions of SPLADE. We would also like to experiment with better ways to perform document expansion, as it is currently the computational bottleneck before training DeepImpact and performing inferences on a new dataset. Ideally, we would like to move away from predicting queries to instead predicting relevant query terms. This will allow us to perform a greater degree of unique term expansion, consequently reducing the issue of vocabulary mismatch further.

### v2

@@ 3.4 Recall at Different Depths
3.5 BEIR Zero-Shot Evaluations
The BEIR (Benchmarking IR) benchmark is a comprehensive evaluation suite that encompasses a wide variety of Information Retrieval (IR) datasets designed to rigorously test the zero-shot effectiveness of retrieval models. To comprehensively evaluate the effectiveness of DeeperImpact  we benchmarked it against SPLADEv3-doc, SPLADEv3, and the large efficient SPLADE variant BT-SPLADE-L [41] on a subset of BEIR datasets using NDCG@10 as the metric. The zero-shot inferences were performed with DeeperImpact model trained on MS MARCO passage dataset. We also used the Llama 2 model—fine-tuned for query prediction on MS MARCO passage dataset—for expanding each of these datasets with relevant queries. The results are displayed in Table 5.
Among the SPLADE variants, SPLADE-v3 consistently achieves the highest NDCG@10 scores across the datasets, demonstrating its state-of-the-art effectiveness. However, it is worth noting that SPLADEv3 incorporates query processing, making direct comparisons less straightforward.
SPLADE-v3-doc, which processes queries as a bag of words without additional query processing, is a more appropriate baseline for comparison with DeeperImpact. In this context, DeeperImpact exhibits competitive effectiveness, often closely aligning with SPLADEv3-doc. Its effectiveness is also fairly close with the large efficient SPLADE variant.
@@ 3.5 BEIR Zero-Shot Evaluations
Table 5: NDCG@10 results across a subset of BEIR datasets for DeeperImpact, BT-SPLADE-L, SPLADE-v3-doc, and SPLADE-v3.
@@ 4 Conclusions and Future Work
In this paper, we presented DeeperImpact  an optimized version of the DeepImpact sparse retrieval model, aimed at improving retrieval effectiveness while maintaining computational efficiency. By substituting DocT5Query with a fine-tuned Llama 2 model for document expansion, leveraging advanced training strategies such as initializing with CoCondenser, employing hard negatives, and integrating distillation techniques, DeeperImpact substantially narrows the effectiveness gap with state-of-the-art but more expensive sparse retrieval models such as SPLADE. Future work could focus on further enhancing the training of DeeperImpact using a combination of training loss objectives and searching for optimal hyperparameters similar to what has been done with recent versions of SPLADE. We would also like to experiment with better ways to perform document expansion, as it is currently the computational bottleneck before training DeepImpact and performing inferences on a new dataset. Ideally, we would like to transition from auto-regressively predicting queries to sampling relevant query terms from a query likelihood distribution—drawing inspiration from the TILDE expansion algorithm. This new approach would not only allow for a greater degree of unique term expansion but also offer significantly enhanced efficiency, thereby further mitigating the issue of vocabulary mismatch.